# URDFormer: A Pipeline for Constructing Articulated Simulation Environments from Real-World Images

### v2

@@ IV Experiments
1. Real-to-sim: First, our system takes an image of the scene and uses a finetuned Grounding DINO [32] (Appendix C-A)to automatically detect the parts of the cabinet, i.e., door, drawer, and handles. URDFormer then generates the corresponding URDF from the RGB image and the predicted bounding boxes using URDFormer, as described in Section II-C.
2. Policy Learning in Sim: The URDF is then imported into a physics simulator (PyBullet [33]) and scaled appropriately using depth measurement. Next, a motion planner, in this case cuRobo [24], generates trajectories that solve a variety of tasks, e.g., closing/opening a drawer/door by utilizing privileged information in simulation. To account for possible prediction inaccuracies of URDFormer (red highlights in Fig 5) and to robustify the trained policy, we apply targeted randomization (TR) that randomizes the scene while maintaining its semantic configurations. In particular, doors, drawers, and handles are randomly replaced with their PartNet [19] equivalents while the size and base of the object are kept fixed. For each door or drawer that was generated in this process, the geometry was randomly replaced with alternate geometry from the same PartNet class, but rescaled to be the appropriate size. Additionally each handle or knob that was generated was similarly replaced with alternate geometry, but also randomly translated in the plane of the door or drawer that it was attached to. Textures are randomized by cropping out the real texture using the bounding boxes, generating variations by prompting Stable Diffusion [18] and fitting it back onto the shape. Finally, the RGB input is augmented by adding standard augmentations such as Gaussian noise and color jitter. After automatically collecting a dataset of successful simulation trajectories using the motion planner, we train a behavior cloning policy network that predicts end-effector poses from point clouds. The network architecture follows M2T2 [25] and predicts 6D end-effector poses from RGB pointclouds given language instructions specifying the task. Appendix E-C provides the full training procedure and architectural details.
@@
(1) OWL-ViT [29]: First, we evaluate against a zero-shot vision-language model baseline. Inspired by VoxPoser [34], given a language instruction (Appendix E-B), we used the same open-vocabulary detector (OWL-ViT [29]) to predict bounding boxes for the prompted parts and handles that are important to the task. After mapping the detection to the observed pointcloud, a motion planner can then generate plans to solve the task directly in the real world.
@@
Real-world Results of Real-to-Sim-Real Training: Running the zero-shot OWL-ViT, we find that the model fails to predict the fine-grained details required to solve the tasks, i.e., ”top middle drawer”, ”right door”, and ”handles”. We visualize qualitative examples of results from OWL-ViT in Appendix E. Without localizing these regions of interest, the motion planning cannot solve the tasks leading to 0% success on all tasks.
@@ IV-B2 Object and Part Detection during Inverse Phase
URDFormer takes both RGB image and bounding boxes of the object parts and predicts URDFs. During the inverse phase, we adopt an off-the-shelf open vocabulary object detector. In particular, we finetune the pretrained GroundingDINO [32] on the generated dataset described in IV-B1. While observing notable improvement on the F1 score, we also found pre-trained network is generally better at detecting out-of-distribution objects and scenes, such as unique handles, or cabinets with special textures. Inspired by recent work Model Soup [37], we combine the finetuned and pretrained GroundingDINO by simply averaging their weights, and observe surprisingly significant improvements that outperform both models. We report both quantitative and qualitative results, as well as details on how we finetune GroundingDINO in Appendix C-A.
@@ IV-B5 Qualitative Results
Figure 7 shows several examples of kitchens generated by URDFormer using internet images. While the model makes some mistakes, it is able to reproduce kitchen environments that largely match the structure of the original images. Figures 8 and 9 show successful and unsuccessful reconstructions of individual objects and scenes respectively. These images also show results when ground-truth boxes are provided. Again, while the model makes some mistakes, it largely captures the overall configuration of the objects and scenes.
@@ IV-B7 Reality Gym
With the ability to cheaply generate an arbitrary number of realistic simulation assets directly from the internet, we introduce RealityGym, a robot learning suite with a collection of realistic simulation assets and scenes created from real world RGB images using URDFormer. We provide an initial set of 300 objects (Cabinets, Ovens, Fridges, wahsers and Dishwashers) and 50 kitchen scenes from internet images, with future work looking to expand this into a bigger dataset. In addition, we also provide 84 meshes of cabinet frames, 20 door meshes, 59 drawers, 440 handles and 116 knobs from PartNet[19]. We can randomly incorporate these meshes into the initial URDFs to generate diverse scenes complete with articulated objects. We define 4 main tasks: (1) Open any articulated parts i.e. top middle drawer (2) Close any articulated parts (3) Fetch objects (4) Collect objects. We automatically generate tasks and their language descriptions, and use a motion planner (Curobo [24]) to complete the tasks. Fig 6 and Fig 7 show examples of robots performing in RealityGym on a variety of generated simulation environments and assets. Details about RealityGym can be found in Appendix D.
@@ IV-D Can URDFormer support different robots and tasks?
To demonstrate that URDFormer supports multi-step tasks and different robots, we train an additional policy for a Stretch robot to place an object in a desk drawer as shown in Figure 12. Figure 13 shows the data generation process for training the mobile robot’s policy. First, a single image of the robot’s environment is passed to a pretrained URDFormer model to produce a scene description of the desk. Then, this scene description is used to generate training data in simulation that can instruct the robot how to accomplish it’s multi-stage objective. When generating this data, we use an inpainting model [18] to reduce the sim2real gap by inpainting the pixels covered by the simulated object overlaid onto the original image. Finally the robot policy, which predicts a per-step affordance map from an initial image, is trained on this new generated data. This policy is implemented as a UNet, which takes the initial observation, as well as the task embedding and predicts multiple object affordance maps. These affordance maps are used with a motion planner to guide the robot at each step. Details on training the policy can be found in Appendix E-C.
@@ B-B Training Details
Appendix C Bounding Box Detection
C-A Fine-tuning Grounding DINO
URDFormer takes images of the object as well as 2D boxes of its parts. However, if we directly apply an off-the-shelf object detector such as GroundingDINO, the detection performance is unsatisfying with an F1 score of 53.4%. Instead, we use the generated dataset described in Appendix A and finetune groundingDINO for detailed bounding box detection such as door, drawer and handle. We evaluate the finetuned groundingDINO on the same dataset and observe an improvement with an F1 score of 66.2%. However, We also observe that compared to the pretrained GroundingDINO, the finetuned GroundingDINO often fails to detect parts that have unique shapes or patterns. Inspired by recent work Model Soup[91], we simply average the pretrained weights and the finetuned weights. This leads to surprising improvement, with an F1 score of 79.7%. We additionally apply post-processing to remove duplicated boxes. Figure 19 shows Model Soup’s influence on the bounding box detection.
@@ C-A Fine-tuning Grounding DINO
Appendix D Reality Gym
D-A Realistic Assets from Images
@@
D-B Targeted Domain Randomization
@@
Appendix E Real-World Experiments
E-A Data Collection in Sim
@@ E-C Policy Training
Appendix F Limitations and Future Work
@@
Appendix G Generalization on Diverse Objects and Scenes

### v3

@@ IV Experiments
1. Real-to-sim: First, our system takes an image of the scene and uses a finetuned Grounding DINO [32] with model soup approach[33] (Detailed in IV-B2) to automatically detect the parts of the cabinet, i.e., door, drawer, and handles. URDFormer then generates the corresponding URDF from the RGB image and the predicted bounding boxes using URDFormer, as described in Section II-C.
2. Policy Learning in Sim: The URDF is then imported into a physics simulator (PyBullet [34]) and scaled appropriately using depth measurement. Next, a motion planner, in this case cuRobo [24], generates trajectories that solve a variety of tasks, e.g., closing/opening a drawer/door by utilizing privileged information in simulation. To account for possible prediction inaccuracies of URDFormer (red highlights in Fig 5) and to robustify the trained policy, we apply targeted randomization (TR) that randomizes the scene while maintaining its semantic configurations. In particular, doors, drawers, and handles are randomly replaced with their PartNet [19] equivalents while the size and base of the object are kept fixed. For each door or drawer that was generated in this process, the geometry was randomly replaced with alternate geometry from the same PartNet class, but rescaled to be the appropriate size. Additionally each handle or knob that was generated was similarly replaced with alternate geometry, but also randomly translated in the plane of the door or drawer that it was attached to. Textures are randomized by cropping out the real texture using the bounding boxes, generating variations by prompting Stable Diffusion [18] and fitting it back onto the shape. Finally, the RGB input is augmented by adding standard augmentations such as Gaussian noise and color jitter. After automatically collecting a dataset of successful simulation trajectories using the motion planner, we train a behavior cloning policy network that predicts end-effector poses from point clouds. The network architecture follows M2T2 [25] and predicts 6D end-effector poses from RGB pointclouds given language instructions specifying the task. Appendix D-C provides the full training procedure and architectural details.
@@
(1) OWL-ViT [29]: First, we evaluate against a zero-shot vision-language model baseline. Inspired by VoxPoser [35], given a language instruction (Appendix D-B), we used the same open-vocabulary detector (OWL-ViT [29]) to predict bounding boxes for the prompted parts and handles that are important to the task. After mapping the detection to the observed pointcloud, a motion planner can then generate plans to solve the task directly in the real world.
@@
Real-world Results of Real-to-Sim-Real Training: Running the zero-shot OWL-ViT, we find that the model fails to predict the fine-grained details required to solve the tasks, i.e., ”top middle drawer”, ”right door”, and ”handles”. We visualize qualitative examples of results from OWL-ViT in Appendix D. Without localizing these regions of interest, the motion planning cannot solve the tasks leading to 0% success on all tasks.
@@ IV-B2 Object and Part Detection during Inverse Phase
URDFormer takes both the RGB image and bounding boxes of the object parts and predicts URDFs. During the inverse phase, we adopt an off-the-shelf open vocabulary object detector GroundingDINO [32].
However, if we directly apply GroundingDINO on detecting parts such as drawers and handles, the detection performance is unsatisfying with an F1 score of 53.4%. Instead, if we use the same generated dataset that was used to train URDFormer (Visualized in Appendix A) and finetune groundingDINO, we observe an improvement with an F1 score of 66.2%. However, We also observe that compared to the pretrained GroundingDINO, the finetuned GroundingDINO often fails to detect parts that have unique shapes or patterns. Inspired by recent work Model Soup[33], we simply average the pretrained weights and the finetuned weights. This leads to surprising improvement, with an F1 score of 79.7%. We additionally apply post-processing to remove duplicated boxes. Figure 8 shows Model Soup’s influence on the bounding box detection.
@@ IV-B5 Qualitative Results
Figure 7 shows several examples of kitchens generated by URDFormer using internet images. While the model makes some mistakes, it is able to reproduce kitchen environments that largely match the structure of the original images. Figures 9 and 10 show successful and unsuccessful reconstructions of individual objects and scenes respectively. These images also show results when ground-truth boxes are provided. Again, while the model makes some mistakes, it largely captures the overall configuration of the objects and scenes.
@@ IV-B7 Reality Gym
With the ability to cheaply generate an arbitrary number of realistic simulation assets directly from the internet, we introduce RealityGym, a robot learning suite with a collection of realistic simulation assets and scenes created from real world RGB images using URDFormer. We provide an initial set of 300 objects (Cabinets, Ovens, Fridges, wahsers and Dishwashers) and 50 kitchen scenes from internet images, with future work looking to expand this into a bigger dataset. In addition, we also provide 84 meshes of cabinet frames, 20 door meshes, 59 drawers, 440 handles and 116 knobs from PartNet[19]. We can randomly incorporate these meshes into the initial URDFs to generate diverse scenes complete with articulated objects. We define 4 main tasks: (1) Open any articulated parts i.e. top middle drawer (2) Close any articulated parts (3) Fetch objects (4) Collect objects. We automatically generate tasks and their language descriptions, and use a motion planner (Curobo [24]) to complete the tasks. Fig 6 and Fig 7 show examples of robots performing in RealityGym on a variety of generated simulation environments and assets. Details about RealityGym can be found in Appendix C.
@@ IV-D Can URDFormer support different robots and tasks?
To demonstrate that URDFormer supports multi-step tasks and different robots, we train an additional policy for a Stretch robot to place an object in a desk drawer as shown in Figure 13. Figure 14 shows the data generation process for training the mobile robot’s policy. First, a single image of the robot’s environment is passed to a pretrained URDFormer model to produce a scene description of the desk. Then, this scene description is used to generate training data in simulation that can instruct the robot how to accomplish it’s multi-stage objective. When generating this data, we use an inpainting model [18] to reduce the sim2real gap by inpainting the pixels covered by the simulated object overlaid onto the original image. Finally the robot policy, which predicts a per-step affordance map from an initial image, is trained on this new generated data. This policy is implemented as a UNet, which takes the initial observation, as well as the task embedding and predicts multiple object affordance maps. These affordance maps are used with a motion planner to guide the robot at each step. Details on training the policy can be found in Appendix D-C.
@@ B-B Training Details
Appendix C Reality Gym
C-A Realistic Assets from Images
@@
C-B Targeted Domain Randomization
@@
Appendix D Real-World Experiments
D-A Data Collection in Sim
@@ D-C Policy Training
Appendix E Limitations and Future Work
@@
Appendix F Generalization on Diverse Objects and Scenes